10.Whichofthefollowingpairsoflinearequationshasuniquesolution,nosolutionorInfinitely
manysolutions?Incasethereisauniquesolution,finditbyusingcrossMultiplicationmethod.
3x-5y=20;6x-10y=40

Grade:10

1 Answers

Aditya Gupta
4 years ago
hello karthik, the given eqns are:
3x – 5y=20......(a)
6x – 10y=40.....(b)
now if we multiply (a) on both sides by 2, we get
2*(3x – 5y)= 2*20
or 6x – 10y= 40, which is precisely the same as (b). hence, (a) and (b) are equivalent to each other.
hence there are infintely many solutions, such as (0, – 4); (5, – 1); (10, 2) etc.
kindly approve :=)